ELECTRONIC SPECTROSCOPY OF THE JET-COOLED ARSENIC DICARBIDE (C_{2}As) FREE RADICAL

\underline{JIE WEI}, FUMIE X. SUNAHORI, ROBERT A. GRIMMINGER AND DENNIS J. CLOUTHIER, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

The linear arsenic dicarbide (C_{2}As) free radical has been identified by observation of the ^{2}Δ - ^{2}Π band system in the 585.0 - 687.3 nm region by laser-induced fluorescence (LIF) and supersonic expansion cooling techniques. The radical was produced in a pulsed electric discharge jet using a precursor mixture of methane (or acetylene), arsenic trichloride vapor and argon. All three excited state vibrational modes have been identified in the LIF spectra of ^{12}C_{2}As and ^{13}C_{2}As. Single vibronic level emission spectra exhibit an isotope-independent ground state spin-orbit splitting of 874 cm^{−1} and activity in the bending and both stretching modes. The measured ground state energy levels have been successfully fitted to a ^{2}Π state Renner-Teller model. High resolution spectra of the 0_{0}^{0} bands of both isotopomers have been recorded and analyzed to obtain the rotational constants and ground and excited state molecular structures. The experimental derived parameters will be compared to ab initio predictions and experimental data available for the analogous C_{2}N and C_{2}P radicals.